A firm has production function f(z) = 2z
1/2 The price of the input is r and the price of
output is p.
a) Calculate the optimal output, input, and profit?
1
Expert's answer
2020-12-04T05:16:08-0500

"f(z) = 2z^{\frac{1}{2}}"

Let:

price of input = r

price of output = p

Profit = Total revenue (TR) – Total cost (TC)

"Profit = p \times 2z^{\frac{1}{2}} – (rz) = 2pz^{\frac{1}{2}} – rz"

To find optimal profit we will differentiate profit function with respect to z.

"\frac{d(profit)}{dz} = 2p \times \frac{1}{2\sqrt{z}} – r = 0 \\

2p = r2\sqrt{z} \\

p = r\sqrt{z} \\

Z^* = (\frac{p}{r})^2"

Optimal input:

"Z^* = \frac{p^2}{r^2} \\

Output = f(z) = 2(Z^*)^{\frac{1}{2}} \\

= 2(\frac{p}{r})^{2 \times \frac{1}{2}} \\

= 2\frac{p}{r} \; (optimal \; output) \\

Profit = 2p(\frac{p}{r})^{2 \times \frac{1}{2}} – r(\frac{p^2}{r^2}) \\

= 2p(\frac{p}{r}) – r \frac{p^2}{r^2} \\

= 2\frac{p^2}{r} - \frac{p^2}{r} \\

= \frac{p^2}{r} \; (optimal \; profit)"